OpenStudy (anonymous):

solve for x.... x^4-x^2-12=0

OpenStudy (anonymous):

Factor

OpenStudy (anonymous):

idk how to factor that??

OpenStudy (anonymous):

(x^2 )(x^2 ) What two number add up to -1 and multiply to -12 ( Hint: One is positive, other is negative)

OpenStudy (anonymous):

im pretty sure that you can only do that with x^2 as the biggest term...

OpenStudy (anonymous):

x^2 = y y^2 - y - 12 = 0 ( y + 3 ) ( y - 4 ) = 0 y = - 3 y = 4 x^2 = y x^2 = 4 x = 2 x^2 = - 3 x = sqrt -3 x = sqrt 3 i i = sqrt -1

OpenStudy (anonymous):

isn't there a different way to do it? i feel like that isn't the traditional way??

OpenStudy (anonymous):

i will be willing to bet that this class does not use complex numbers

OpenStudy (anonymous):

It does not need to be only x^2 to factor

OpenStudy (anonymous):

You can factor any exponent like this

OpenStudy (anonymous):

how...can you just show it?

OpenStudy (anonymous):

(x^2-4)(x^2+3)

OpenStudy (anonymous):

not really sure how you got there??

OpenStudy (anonymous):

x^2=4 x=+ or - 2

OpenStudy (anonymous):

and that they want you to say "let \[y=x^2\] and solve \[y^2-y-12=0\] \[(y-4)(y+3)=0\] \[y=4\] or \[y=-3\] \[x^2=4\] or \[x^2=-3\] second one has no real number solutions. first one has solutions \[x=2\] or \[x=-2\]